OpenStudy (anonymous):

Simplify: 6y + 5z + 8y – 3z A. 11y + 5z B. 14y + 2z C. 14y + 8z D. 16yz

OpenStudy (sleepyhead314):

add like terms you can only add together terms that have the same variables attached to them for example 6y + 5z cannot be added but 6y + 8y can do you know what that would equal?

OpenStudy (anonymous):

@sleepyhead314 no i don't:/

OpenStudy (sleepyhead314):

for example: 4y + 7y = 11y like how 4+7=11 but it's just with the y variables

OpenStudy (anonymous):

ohh lol i c

OpenStudy (anonymous):

so i just add the numbers as an a equation and but the letters next to it? @sleepyhead314

OpenStudy (sleepyhead314):

yes ^_^ so what do you think 6y + 8y = ?

OpenStudy (anonymous):

15 ^_^

OpenStudy (sleepyhead314):

close :) don't forget the y so 6y + 8y = 15y get it now?

OpenStudy (anonymous):

15y

OpenStudy (sleepyhead314):

^ yes :))

OpenStudy (sleepyhead314):

then it's the same idea with the z for your question

OpenStudy (anonymous):

ao to simplify it what do i do? @sleepyhead314

OpenStudy (sleepyhead314):

in this case, simplifying just means to add like terms :)

OpenStudy (anonymous):

and so its 5+-3?

OpenStudy (sleepyhead314):

yep :D

OpenStudy (anonymous):

so its 6+8 + 5-3?

OpenStudy (anonymous):

OHHH I THINK I GET IT @sleepyhead314

OpenStudy (sleepyhead314):

^_^ just don't forget the variables :)

OpenStudy (anonymous):

is it 17yz? O.o @sleepyhead314

OpenStudy (sleepyhead314):

not quite, remember when I said that y and z cannot be added? ;) so just 15y + 2z

OpenStudy (sleepyhead314):

whoops I meant *14y + 2z

OpenStudy (sleepyhead314):

sorry, I've gtg to school, I hope I was able to help you, ttyl

OpenStudy (anonymous):

SO ITS B

OpenStudy (sleepyhead314):

yes

OpenStudy (anonymous):

thanks so much @sleepyhead314 saved my life<3 hope to ask you more question again :D